OpenStudy (anonymous):

can some one please help me do this ∫4x^2(2x^3-7)^2 dx step by step?

OpenStudy (anonymous):

Im supposed to find the indefinite integral

OpenStudy (campbell_st):

it can be rewritten \[4\int\limits x^2(2x^3 -7)^2 dx\] use substitution let \[u = 2x^3 - 7\] then \[\frac{du}{dx} = 6x^2\] so \[du = 6x^2 dx\] or \[x^2 dx = \frac{1}{6} du\] so using the substitution \[4timesfrac{1}{6} \int\limits u^2 du\]

OpenStudy (campbell_st):

oops last line should read\[4 \times \frac{1}{6} \int\limits u^2 du\]

OpenStudy (anonymous):

thats not the answer in my book....

OpenStudy (campbell_st):

which gives \[\frac{2}{3} \times \frac{1}{3} u^3 + c\] and \[u = 2x^3 - 7 \] gives \[\frac{2}{9} (2x^3 - 7)^3 + C\]

OpenStudy (campbell_st):

hope thats the answer

OpenStudy (anonymous):

haha yeah it is.

OpenStudy (campbell_st):

so its all about the correct substitution... and working from there

OpenStudy (campbell_st):

to check your solution you can differentiate

OpenStudy (anonymous):

ummm can you please explain how to get to step 2? haha sorry i'm retarded.

OpenStudy (campbell_st):

is that the derivative..?

OpenStudy (anonymous):

no after you substitute u.

OpenStudy (campbell_st):

is it \[4 \times \frac{1}{6} \int\limits u^2 du\]

OpenStudy (campbell_st):

perhaps if you draw it it may be easier

OpenStudy (anonymous):

no the \[du/dx=6x^2\] part

OpenStudy (anonymous):

how did you get that?

OpenStudy (campbell_st):

ok.... look at the integral in the rewritten form \[4 \int\limits x^2(2x^3 - 7) dx\] I can group things \[4 \int\limits (2x^3 - 7) x^2dx\] and if \[u = 2x^3 - 7\] I need to find the derivative of u and get it to match x^2 dx in the integral

OpenStudy (campbell_st):

so \[\frac{du}{dx} = 6x^2\] manipulating the derivative \[du = 6x^2 dx\] but I only want x^2 dx... so I need to divide both sides by 6 hence \[\frac{1}{6} du = x^2 dx\] so I really make 2 substitutions 1 for \[u = 2x^3 - 7\] and one for \[\frac{1}{6} du = x^2 dx\]

OpenStudy (anonymous):

HOW DID YOU DO THAT FIRST PART?

OpenStudy (campbell_st):

if it was \[y = 2x^3 - 7\] can you differentiate it with respect to y...?

OpenStudy (anonymous):

NO CAN YOU HELP ME DO THAT

OpenStudy (campbell_st):

ok.... then is is a big issue.... to be able to do integral calculus you need to know the opposite.. differential calculus... and Its a huge topic... for this question the theory is \[y = x^n\] then \[\frac{dy}{dx} = nx^{n-1}\]

OpenStudy (anonymous):

4∫x2(2x3−7)2dx 4∫x2(4x^6-28x^3+49)dx 4∫4x^8-28x^5+49x^2)dx 4[4∫x^8dx-28∫x^5dx+49∫x^2dx] ∫x^n=x^n+1/n+1 4[4(x^9/9-28x^6/6+49x^3/3] 16x^9/9-4*28x^6/6+4*49x^3/3

OpenStudy (anonymous):

Caitlin1204 understood